### v6

@@ 5 Conclusions
In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise in the presence of non-smooth initial data. For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{\alpha/p}+h^{2/p}t^{-1/p})$ with non-smooth initial data. Meanwhile, under the presence of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the norm of $L^{p}(\Omega\times(0,T);L^{q})$, alongside a pathwise uniform convergence rate of $O(h^{2-\epsilon})$. Significantly, these findings are inclusive of diverse spatial $L^{q}$-norms, enhancing their adaptability and universality. The derived theoretical results set a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise.

### v7

@@ 5 Conclusions
In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, particularly in the presence of non-smooth initial data. For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{\alpha/p}+h^{2/p}t^{-1/p})$ with non-smooth initial data. Meanwhile, under the condition of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the $L^{p}(\Omega\times(0,T);L^{q})$ norm, along with a pathwise uniform convergence rate of $O(h^{2-\epsilon})$. Significantly, these findings cover a variety of spatial $L^{q}$-norms, enhancing their adaptability and universality. The theoretical results derived here lay a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, especially within the context of general spatial $L^{q}$-norms.